A coin was tossed 10 times and the outcomes noted as H,T,H,H,T,T,T,T,H,T where H is Head and T is Tail. Find the estimated probability of a tail.
1
Expert's answer
2012-03-13T09:57:54-0400
We'll use classical definition of probability
Due to it probability of event A P(A)={number of favourable events}/{total number of events}=n/N
In opur case event A-getting a tail.
N=10
n=6 (number of letters T in the list)
Therefore p(A)=6/10=3/5